jagr2713 (jagr2713):

Trivia Question! #2 (No direct answers & Explain)

jagr2713 (jagr2713):

Solve (143(5) + 67)5 ___________ 2

OpenStudy (max1409):

3910

jagr2713 (jagr2713):

@Nnesha @geerky42 @EclipsedStar @iGreen @Whitemonsterbunny17

jagr2713 (jagr2713):

Nope @max1409 but nice try :D

eclipsedstar (eclipsedstar):

\(\color{blue}{\text{Originally Posted by}}\) @jagr2713 Solve (143(5) + 67)5 ___________ 2 \(\color{blue}{\text{End of Quote}}\) What's with that lone 2 sitting over there? o.o

jagr2713 (jagr2713):

Hmm you tell me

HanAkoSolo (jamierox4ev3r):

I just noticed that o-o

OpenStudy (anonymous):

Multiply 143 and 5. Add 67 to the result and then multiply by 5. 3910

jagr2713 (jagr2713):

Nope nice try thou @robtobey

jagr2713 (jagr2713):

@zepdrix

eclipsedstar (eclipsedstar):

Order of operations doesn't make sense here though

zepdrix (zepdrix):

So it's some type of math question riddle I guess? :) Hmmm lol

OpenStudy (anonymous):

........

jagr2713 (jagr2713):

Yea @zepdrix :D

HanAkoSolo (jamierox4ev3r):

(143(5)+67)5 ________ 2 143*5=715 715+67=782 782*5=3910 3910\(\div\)2=1955 is it 1955? O-o

jagr2713 (jagr2713):

OMG @Jamierox4ev3r for 2 :DDD

eclipsedstar (eclipsedstar):

Yeah that's it

jagr2713 (jagr2713):

We got a math wizz in here :D

HanAkoSolo (jamierox4ev3r):

LOL seriously?! x'D I tried to go along with the parentheses and add multiply accordingly i had NO clue what ___________ meant , so i guessed :P

jagr2713 (jagr2713):

I did the 2 like that b.c people dont generally look and just want to finish first :D

eclipsedstar (eclipsedstar):

Pretty confusing trick there...lol

HanAkoSolo (jamierox4ev3r):

it it wasn't for you my dear musketeer @EclipsedStar , i would have never noticed the two :oooo

eclipsedstar (eclipsedstar):

I saw that 2 and I was like "ummm what's it doing there"

HanAkoSolo (jamierox4ev3r):

XD

eclipsedstar (eclipsedstar):

x'D

jagr2713 (jagr2713):

Yea some people dont really look lol

jagr2713 (jagr2713):

Again great job @Jamierox4ev3r and thanks for trying guys

Nnesha (nnesha):

@jagr2713 surprise!!!! 4 u http://openstudy.com/users/tkhunny#/updates/5575912ee4b0639420678313 :3

Nnesha (nnesha):

today 12 hours ago^

jagr2713 (jagr2713):

lol didnt see that

Nnesha (nnesha):

now jgr like http://prntscr.com/7escag :P